OpenStudy (anonymous):

Management of the Sycamore Steel Company wishes to determine if there is any difference in performance between the day and night shift workers. They randomly sample a group of 40 day-shift workers and found a mean output of 62 parts per hour with a standard deviation of 9 parts per hour. A random sample of 30 night shift workers found an output of 67 parts per hour and a standard deviation of 8 parts per hour. At the 0.05 level of significance is there a difference between the day and night shift works?

OpenStudy (anonymous):

@satellite73

OpenStudy (dumbcow):

you weren't able to get the answer ?

OpenStudy (anonymous):

The equation you provided was the wrong equation sorry. :-(

OpenStudy (dumbcow):

oh im sorry, i assumed this is comparing 2 sample means

OpenStudy (anonymous):

It is but the means are different.

OpenStudy (anonymous):

Sorry your formula was correct after all I'm used to very different notation that's all. :-)

OpenStudy (dumbcow):

oh ok , i was going to say it does also depend on whether the population variance is equal or not i think you should assume equal variance here

OpenStudy (anonymous):

So, here's what I have so far, Sp=sqrt(73.75) T=1.6676

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

yes thats what i get for Sp but i get a different test statistic \[T = \frac{67-62}{S_p \sqrt{\frac{1}{40} + \frac{1}{30}}} = \frac{5}{\sqrt{73.75} \sqrt{\frac{7}{120}}} = 2.41\] critical value for df =68, alpha = .05 is about 2 http://www.sjsu.edu/faculty/gerstman/StatPrimer/t-table.pdf Therefore you would reject null hypothesis and conclude there is significant difference in performance between day and night shifts @rosho